# M3LEO: A Multi-Modal, Multi-Label Earth Observation Dataset Integrating Interferometric SAR and RGB Data

### v1

@@ 3.2 Labelled Datasets & Digital Elevation Models
Land cover classification labels (semantic segmentation) were obtained from the ESA World Cover product (ESAWC) [77], as available in Google Earth Engine [^4] https://developers.google.com/earth-engine/datasets/catalog/ESA_WorldCover_v200. The resolution of ESAWC is 10m/pixel, and comprises 11 classes (See Appendix C). Data is provided for 2020.
@@
A map of above ground biomass (AGB) in Mg ha-1, derived from the ESA Climate Change Initiative’s Biomass product [^5] https://gee-community-catalog.org/projects/cci_agb/[78] is provided for 2020 at a resolution of 90m/pixel.
@@
Tree cover, non-tree cover and non-vegetated (bare) percentage labels derived from the Terra MODIS Vegetation Continuous Fields product (MODISVEG) [^6] https://developers.google.com/earth-engine/datasets/catalog/MODIS_006_MOD44B[79], are provided at a resolution of 250m/pixel. Our dataset includes yearly maps for 2016-2020.
@@
Maps of built surface area (m2/pixel), derived from the Copernicus Global Human Settlement Built Surface (GHSBUILTS) product [^7] https://human-settlement.emergency.copernicus.eu/ghs_buS2023.php[80] are provided at 100m/pixel for 2020.
@@
Digital Elevation Maps, derived from the NASA Shuttle Radar Topopgraphic Mission (SRTM) [^8] https://developers.google.com/earth-engine/datasets/catalog/CGIAR_SRTM90_V4[81][82] are provided at a resolution of 90m/pixel. This data was measured in 2000, but we emphasise that terrain height changes relatively little at this resolution compared to other products such as ESAWC or GHSBUILTS.
@@ 3.3 Data coverage
Our dataset covers 6 distinct geographic areas of interest (AOIs), comprising the contiguous United States (CONUS), Europe, the Middle East, Pakistan and India (PAKIN), China, and South America. A visualisation is provided in Figure 1. We chose for our data to span only the area covered by the ARIA S1 GUNW dataset from which we obtained interferometric data. Many components of our dataset, such as Sentinel-1 GRD, have far greater coverage. We opt not to fully include these in the initial release to reduce storage requirements, as we aim to encourage the use of the full diversity of SAR datatypes. Processing such data for usage in our framework is technically straightforward using the provided tools, should the additional coverage be required. Note that we include data for Europe despite the absence of GUNW for this region, due to interest from data providers operating in Europe.
A total of 1,048,827 unique geographic tiles were generated, covering an area of $2.11\times 10^{7}$km2. A breakdown by AOI can be seen in Table 1. Tiling is uniform across all AOIs and datasets - the chips provided for each dataset cover exactly the same geographical areas. It is important to note that not all component datasets or specific parameterizations, such as date-pair ranges, were available for every geographic tile. However, the availability is still consistently high. Our framework addresses these discrepancies by training and testing selectively on tiles for which all desired datasets are available. See Appendix A for a full breakdown by dataset and AOI. We provide a reduced version of our dataset, M3LEO-miniset, spanning 5000 tiles per AOI for rapid model iteration and use in tutorials.
@@
We used geographic banding to define training, validation and test splits. Bands were split into training, validation and test sets at a ratio of 60:20:20, and can be seen visually in Figure 1. This method reduces distribution shifts and data leakage compared to single-band splits and fully randomized assignments, respectively. See Appendix A for details.
@@ 3.4 Framework
To process datasets available via Google Earth Engine (GEE) (S2RGB, S1GRD, SRTM, ESAWC, MODISVEG), we introduce geetiles [^9] https://github.com/rramosp/geetiles. This tool extracts and tiles data from GEE as per definition files and configurations provided in the geetiles repository. The remaining datasets (GSSIC, GUNW, AGB, GHSBUILTS) were extracted using sartiles [^10] https://github.com/rramosp/sartiles, which contains specific code to download and tile each of GUNW and GSSIC, as well as the facility to tile general GeoTIFF files, such as those provided for AGB and GHSBUILTS, for integration into M3LEO. Both geetiles and sartiles were developed alongside M3LEO, and are provided such that users are able to seamlessly integrate any data available via Google Earth Engine (or as a GeoTIFF file) with our framework.
Deep Learning Pipeline
We accompany our dataset with a modular PyTorch Lightning [83] framework parameterised using Hydra [28]. We provide PyTorch Lightning datasets for each component of M3LEO as well as the facility to use multiple datasets flexibly as input data or targets. We provide PyTorch Lightning modules defining the UNet-style models used in this work. We also provide additional modules defining self-supervised approaches applied successfully to SAR data in previous works (MAE [63], CLIP [65], DINO [66][67]). Integrating custom models with M3LEO is straightforward, requiring the addition of a single file.
@@
3.5 Baseline Evaluations
Data
To evaluate the efficacy of our data and framework for Earth observation tasks, we trained small UNet-style models on a subset of the input and labelled datasets. S1GRD, GSSIC coherence and S2RGB were used as input datasets. For S1GRD, we trained models separately using the VV and VH bands only, and with the VV and VH bands stacked at the input to the model. For all S1GRD models, four seasonal summaries were used for each band, resulting in four input channels for the VV and VH models, and eight channels for the model using both VV and VH. For GSSIC coherence, the coherence band was used with a single date pair of delta 36 days taken for each season, resulting in four input channels. For the S2RGB models, we used the red, green and blue channels with one input channel from each of the months of March, June, September and December, totalling 12 input channels. We additionally trained models using both S2RGB in combination with each of the other datasets separately, stacking the bands at the input to the model. All input data was taken from 2020. We resized low resolution input datasets to $448\times 448$ so that all model inputs had a resolution of 10m/pixel, ensuring that the each layer in our model operates on the same spatial scale for all datasets.
We excluded GUNW from use in our baseline experiments to avoid introducing the mixed availability of GUNW chips as a confounding factor. SRTM data was excluded as an input due to budgetary constraints, as low-resolution terrain data likely has less predictive power than the selected datasets. We encourage users to find more nuanced uses for these datasets. We also excluded the MODISVEG task as sensible use of resources, as results may be heavily correlated with the AGB task.
The ESAWC, AGB and GHSBUILTS labelled datasets were selected as targets in our baseline experiments. ESAWC is formulated as semantic segmentation spanning 11 land use classes, for which segmentation accuracy is reported as mean intersection-over-union (mIoU). ESAWC data was used at the original resolution of $448\times 448$. Both AGB and GHSBUILTS are formulated as regression tasks (per-pixel). Results are reported using root mean square error (RMSE), in Mg ha-1 for AGB and in m2 built surface for GHSBUILTS. We resized labels for AGB and GHSBUILTS to a fixed size of $45\times 55$, to account for minor differences in dimension from chip-to-chip. We note that this means our pixels may not span exactly 1 hectare and therefore that results for AGB measured in Mg ha-1 are a heuristic rather than an absolute measure of biomass.
We excluded data from Europe in our baseline experiments due to the absence of GUNW data in this region. This constraint was applied despite having not used GUNW in these baselines, as these experiments were completed prior to the exclusion of GUNW and repeating them was prohibitively expensive. For all experiments, we used a randomly selected $10\%$ of the training and validation sets to decrease computational cost. The same selection was used across all experiments. We used the entire test set to compute the final metrics.
Models
We used a UNet-style architecture for all baseline experiments. We follow the architecture of [84], with two major changes - halving the number of channels for all layers and replacing the up-convolutions with bilinear upsampling. We chose to make these adjustments to reduce computational cost. Additionally, for the AGB and GHSBUILTS regression tasks, we applied average pooling to the single-channel $448\times 448$ UNet output to achieve an output dimension of $45\times 55$. We motivate our choice of model with the observation that UNet style architectures have achieved great success in Earth observation tasks using both RGB and SAR data [85][86][87][88][88]. We further motivate this choice with the observation that vision transformer-based models may be more reliant on data augmentation or pretraining to achieve high performance compared to CNN-based models [89]. Selecting augmentations for SAR data is challenging - common choices such as rotation or flipping may introduce invariance to information specific to different instrument polarisations or orbital direction, for example. We therefore chose not to apply data augmentations in our experiments. For further details on training and hyperparameters, see Appendix B.
3.6 Results
Results for all tasks are reported in Table 2. For the experiments using a single data source as input, S1GRD using both polarisations (VV+VH) achieved the best result for the ESAWC (MIoU: 0.4185) and AGB (RMSE: 27.467 Mg ha-1) tasks. S2RGB achieved the best result for GHSBUILTS (RMSE: 131.968 m2). GSSIC coherence produced the worst result in all three cases (ESAWC MIoU: 0.2906, AGB RMSE: 39.314 Mg ha-1, GHSBUILTS RMSE: 131.968 m2). The best results for experiments using multiple data sources as input were achieved by fusion of S1GRD and S2RGB in all cases (ESAWC MIoU: 0.4634, AGB RMSE: 25.137 Mg ha-1, GHSBUILTS RMSE: 124.852 m2), with significant improvements over either S1GRD or S2RGB individually. Fusion of S2RGB with GSSIC improved results marginally (ESAWC MIoU: 0.4198, AGB RMSE: 28.550 Mg ha-1, GHSBUILTS RMSE: 131.300 m2) compared to either individual data source individually in all cases.
Table 2: Baseline Evaluation Results for ESAWC, AGB, and GHSBUILTS tasks using our UNet models, outlined in Section 3.5.
Input Dataset Bands ESAWC AGB GHSBUILTS
MIoU RMSE (Mg ha-1) RMSE (m2 built)
S1GRD VV 0.3976 28.573 152.259
S1GRD VH 0.3787 30.333 152.847
S1GRD VV+VH 0.4185 27.467 141.719
GSSIC Coherence 0.2906 39.314 196.242
S2RGB RGB 0.4094 28.689 131.968
S2RGB+S1GRD RGB+VV+VH 0.4634 25.137 124.852
S2RGB+GSSIC RGB+Coherence 0.4198 28.550 131.300
4 Discussion
In all experiments, S1GRD combining VV and VH polarisations performed similarly to S2RGB. As we chose not to repeat our experiments to obtain error bars due to computational cost, and given the relatively small gap in performance, we do not claim conclusively that either S1GRD or S2RGB achieves greater performance for any particular task. However, given the sizes of the training (63×$\times 10^{3}$ tiles; 10% of the full training set) and test (105×$\times 10^{3}$ tiles) sets used, we are able to make preliminary suggestions about our data. The inclusion of both polarisations for S1GRD increased performance uniformly compared to either individually. Unlike optical data, reflected SAR pulses are polarised differently depending on the terrain, and each measured polarisation contains unique information about the geometry of the imaged area.
In all cases, performance was significantly improved by using both S1GRD and S2RGB data in fusion, despite giving similar performances individually. We suggest that SAR data contains information orthogonal to that found in RGB imagery. We used RGB channels only, so cannot confirm whether including additional S2RGB bands such as Short-Wave Infrared (SWIR) would reduce this effect.
Baseline experiments using GSSIC as the sole input data source performed significantly worse than for other input data types. This is likely explained in large part by the difference in resolution between GSSIC (90m, upsampled to 10m), S1GRD (10m) and S2RGB (10m). Performance improved slightly in all cases when GSSIC was included alongside in fusion with S2RGB. A nuanced approach is required for including coherence and interferometry, rather than naive direct input. Users may wish to use these data sources in self-supervised learning schemes - for example, pretraining models by constructing coherence or interferometry data from polarimetry pairs, or in a contrastive scheme alongside polarimetry data. Some work has been successful in using this type of pretraining [65]. Motivated by the success of polarimetry and coherence data in non-ML tasks [17][18][19][20][15][13][14][16], we suggest that these data types generated from date-paired SAR acquisitions are likely to show stronger performance on change detection-type tasks, which we did not demonstrate here. We aim to include such task data in a future update to M3LEO. Although the same information found in coherence and polarimetry data could hypothetically be learned directly from paired acquisitions (assuming phase information, which is not provided via Google Earth Engine), the use of pre-computed coherence and interferometric data is likely to be helpful in low-data regimes. Pretraining models for few-shot change-detection tasks using coherence and interferometry may prove to be fruitful.
@@ 5 Conclusions and Future Work
In this work, we introduced M3LEO, a multi-continental Earth observation dataset including a comprehensive set of SAR data, alongside digital elevation models, RGB data and a suite of downstream tasks. To our knowledge, this is the only work to date including SAR datatypes beyond polarimetry in an ML-readable format. We aim to update M3LEO with further component datasets in the future.
We additionally provide a modular PyTorch Lightning framework to enable the application of deep learning and encourage the uptake of these datatypes. We provide additional tools, geetiles and sartiles, to enable the integration of any data available in Google Earth Engine with our framework.
We conducted a small set of experiments to validate the utility of our data and framework. We found that polarimetric SAR amplitude data was competitive with RGB imagery. Results were significantly improved when polarimetric SAR amplitude was used in fusion with RGB data, suggesting that there is useful information found in SAR data in addition to that which can be found in RGB imagery. The naive use of an additional SAR datatype - coherence - did not give comparable results, but this is likely confounded by resolution differences. Motivated by the success of this datatype in other applications [13][14][15][16] we suggest that these data are still useful, but must be applied in a more nuanced way. One such example might include self-supervised learning schemes, where coherence data has already been used successfully [65][67].
@@ 6 Acknowledgements
      Did you describe the limitations of your work? [Yes] See Appendix D.
@@
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix B
@@
      Did you mention the license of the assets? [Yes] See Appendix E.
@@ A.1 Chip counts
[h] Summary of component datasets in M3LEO, including number of chips and dataset size for each region.
@@
    Data undergoing preparation and will be in place prior to camera-ready version.
@@ A.2 Banding
Appendix B Training Hyperparameters

### v2

@@ 3.2 Auxiliary Datasets
[…]
@@
A map of above ground biomass (AGB) in Mg ha-1, derived from the ESA Climate Change Initiative’s Biomass product555gee-community-catalog.org/projects/cci_agb/ [92] is provided for 2020 at a resolution of 90 m/pixel. The relative error of this data is $20\%$ for areas with a measured biomass exceeding $50$ Mg ha-1 [93]
@@
[…]
@@
[…]
@@
Digital Elevation Maps, derived from the NASA Shuttle Radar Topopgraphic Mission (SRTM)888developers.google.com/earth-engine/datasets/catalog/CGIAR_SRTM90_V4[97, 98] are provided at a resolution of 90 m/pixel. This data was measured in 2000, but we emphasise that terrain height changes relatively little at this resolution compared to other products such as ESAWC or GHSBUILTS. The RMSE of the SRTM data was originally reported as 16 m [99] although it has been measured as more accurate in some regions [100].
@@ 3.3 Data coverage
Our dataset covers six distinct geographic areas of interest (AOIs): the contiguous United States (CONUS), Europe, the Middle East, Pakistan and India (PAKIN), China, and South America. A visualisation is provided in Figure 1. We limit coverage to these regions for reasons relating to the acquisition parameters used by Sentinel-1.
Firstly, Sentinel-1 operates with different acquisition modes depending on the region. We choose to only include areas where Sentinel-1 uses the Interferometric Wide (IW) swath acquisition mode. In polar regions, Sentinel-1 employs the Extra Wide (EW) swath acquisition mode, which introduces systematic differences in how terrain is illuminated — such as the azimuth steering angle of the radar emitter being 0.8° for EW and 0.6° for IW acquisition. The polarisations used in polar regions are reversed (HH, HV vs. VV, VH) compared to other areas. We provide IW data with both VV and VH polarisations in the initial release of this dataset.
Secondly, much of the terrestrial surface of the Earth is only covered by a single direction of satellite travel. This includes much of North America, Africa, continental Asia, Oceania and the Amazon rainforest. Unlike passively illuminated data such as Sentinel-2, SAR actively illuminates terrain with a radar emitter aimed laterally from one side of the satellite. Orbital direction therefore systematically determines whether terrain is illuminated from an easterly or westerly direction.
By focusing on regions with consistent acquisition parameters, we aim to provide a dataset that is more uniform and suitable for training models without introducing additional complexities. While including other regions might reduce geographic bias, it is not straightforward to address the potential systematic biases introduced by varying acquisition modes and polarisations, or the systematic lack of directional coverage in some regions. These issues warrant a detailed treatment beyond the scope of this work, and we refer readers to the Copernicus Wiki for a full set of details on Sentinel-1 coverage999sentiwiki.copernicus.eu/web/s1-mission (particularly Figures 20, 21). Processing additional coverage is technically straightforward using the provided tools, should it be required. We include data for Europe despite its absence of GUNW coverage due to interest from data providers operating in Europe.
A total of 1,048,827 unique geographic tiles were generated, covering an area of $2.11\times 10^{7}$ km2. A breakdown by AOI can be seen in Table 1. Tiling is uniform across all AOIs and datasets—the chips provided for each dataset cover exactly the same geographical areas. It is important to note that not all component datasets or specific parameterisations, such as date-pair ranges, were available for every geographic tile; however, the availability is still consistently high. See Appendix A for a full breakdown by dataset and AOI. We provide a reduced version of our dataset, M3LEO-miniset, spanning 5,000 tiles per AOI for rapid model iteration and use in tutorials.
@@
We provide use geographic bands to define training, validation and test splits for use in the explorations foudn in this work, following [65]. Bands were split into training, validation and test sets at a ratio of 60:20:20, and can be seen visually in Figure 1. This method reduces distribution shifts and data leakage compared to single-band splits and fully randomized assignments, respectively. See Appendix A for details. Users should define train-test splits appropriate for their individual applications if their needs differ.
@@ 3.4 Framework
To process datasets available via Google Earth Engine (GEE) (S2RGB, S1GRD, SRTM, ESAWC, MODISVEG), we introduce geetiles101010github.com/rramosp/geetiles. This tool extracts and tiles data from GEE as per definition files and configurations provided in the geetiles repository. The remaining datasets (GSSIC, GUNW, AGB, GHSBUILTS) were extracted using sartiles111111github.com/rramosp/sartiles, which contains specific code to download and tile each of GUNW and GSSIC, as well as the facility to tile general GeoTIFF files, such as those provided for AGB and GHSBUILTS, for integration with M3LEO. Both geetiles and sartiles were developed alongside M3LEO, and are provided such that users are able to seamlessly integrate any data available via Google Earth Engine (or as a GeoTIFF file) with our framework.
Pipeline
We accompany our dataset with a modular PyTorch Lightning [101] framework parameterised using Hydra [29]. We provide PyTorch Lightning datasets for each component of M3LEO. We also provide additional modules defining self-supervised approaches applied successfully to M3LEO in previous works (MAE [65], CLIP [67], DINO [68, 69]). Integrating custom models with M3LEO is straightforward, requiring the addition of a single file.
@@
4 Analysis
We provide five auxiliary datasets (ESAWC, AGB, MODISVEG, GHSBUILTS, SRTM) in addition to SAR and optical satellite data acquisitions. Although these data could be used as labelled tasks (see Appendix D, and also [65, 67, 69, 68] for analysis of this type on M3LEO), they could equally be considered as providing information on important surface properties that would be non-trivial to derive directly from satellite data.
We compare the shift in the marginal distribution of these terrain properties $y$, $p(y)$ in Figure 3 for four of these auxiliary datasets. We computed the $L_{1}$ distance between the normalised discrete ESAWC distributions and the Wasserstein distance between the continuous GHSBUILTS, MODISVEG and SRTM distributions, across AOIs. The marginal distributions $p(y)$ for each auxiliary dataset are shown in Appendix B.
We also provide an early exploration of the ‘appearance shift’ of features between AOIs for S1GRD — the change in the distribution of embeddings $x$ for a SAR tile with known terrain properties $y$, $p(x|y)$. To generate embeddings, we trained a masked autoencoder on S1GRD polarimetry from all six AOIs, following previous work on M3LEO [65]. Training hyperparameters can be seen in Appendix C. We applied max-pooling along the sequence dimension at the output of the ViT encoder and further reduced the dimension to 2 using UMAP [102]. We computed the expected Wasserstein distance between the conditional distributions $p(x|y)$ with respect to the distribution $p(y)$ on the test set, across AOIs and show the results in Figure 4. For ESAWC, we applied principal component analysis and conditioned on the first 3 principal components with 10 evenly spaced bins per dimension. On the remaining variables, we used 100 evenly spaced bins. We also display the sliced Wasserstein distance between the marginal embedding distributions $p(x)$ in the leftmost matrix of Figure 4. We plot these reduced embeddings, coloured according to each of the terrain properties, in Figure 5. A checkpoint for the MAE model is available in the data repository.
Figure 3: Distribution shifts. Distribution shifts between terrain properties described by auxiliary datasets. Distribution shift for the ESAWC categorical data was quantified using L1 distance and continuous data using Wasserstein distance.
Figure 4: Covariance and appearance shift. (Leftmost) The Sliced Wasserstein Distances (SWD) between reduced train and test set embeddings across AOIs, generated using a masked autoencoder. (Right four) The expected value of same metric conditioned on each of four terrain properties. The expectation was computed with respect to the distribution of the terrain property on the test set. We conditioned on the first three principal components of the ESAWC distribution.
Figure 5: Embedding scatter plots. Scatter plots of 2D embeddings, reduced using UMAP, coloured according to different auxiliary datasets.
We observed that there was significant covariate shift in the distribution of the embeddings produced by the masked autoencoder, $p(x)$, across AOIs (Figure 4, leftmost matrix). Given that there was also significant shift in terrain properties described by the auxiliary datasets between AOIs (Figure 3), this is not immediately surprising. Another factor that must be considered, however, is the shift in the embeddings produced by the masked autoencoder for tiles with similar terrain properties $y$, $p(x|y)$, across AOIs (appearance shift). It can be seen in the four right-hand matrices of Figure 4 that although there is some reduction in the most extreme cases, the expected value of sliced Wasserstein distance conditioned on similar terrain properties is usually not substantially lower than the unconditioned case. Explained from an Earth observation perspective - the masked autoencoder does not extract overly similar features for two tiles with, for example, very high vegetation when those tiles are taken from different geographic regions. This effect can be seen visually in Figure 5 - embeddings with high values for particular labels do not cluster obviously across different AOIs. In contrast, previous work applying DINO-based self-supervision to M3LEO [69] found that embeddings with similar labels clustered tightly in the embedding space across AOIs. Despite this observation, previous work applying MAE-based pretraining to M3LEO has shown strong generalisation to novel AOIs [65].
We did not explore the use of GSSIC coherence data here, as this is substantially technically challenging and requires detailed treatment. Given the lower resolution of this data, it may not be productive to use as a naive input to deep learning models. A small set of experiments evidencing this is provided in Appendix D. We point readers to previous work using coherence data from M3LEO productively in a self supervised setting - in contrastive learning [67] and in knowledge distillation-based approaches [69]. We note that the provision of GUNW interferogram data without reference to specific events such as floods or fires is unusual compared to other datasets [76, 77, 78]. Although we aim to include event-based datasets in a future update, this large-scale dataset of interferograms is still highly desirable in self-supervised schemes.
A number of unknowns remain regarding domain shift in M3LEO. Many auxiliary datasets, such as ESAWC, do not exist for 2018 and 2019 so it is difficult to provide a substantial exploration of temporal shifts, although we provide S1GRD and S2SRM for three years. It is unclear whether features learned from encoders trained on different polarizations or orbital directions are comparable, although D contains a limited set of experiments on the value-add of different polarisations.
@@ 5 Conclusions and Future Work
In this work, we introduced M3LEO, a multi-continental Earth observation dataset including a comprehensive set of SAR data, alongside digital elevation models, RGB data and a suite of downstream tasks. To the best our knowledge, this is the largest ML-readable polSAR dataset by total number of tiles and geographic coverage, and the largest inSAR dataset by the same metrics. We additionally provide a modular PyTorch Lightning framework to enable the application of deep learning and encourage the uptake of these datatypes. We provide additional tools, geetiles and sartiles, to enable the integration of any data available in Google Earth Engine with our framework.
We trained an MAE-based model on polSAR data and conducted a small exploration on the appearance shift of features corresponding to similar labels across AOIs. Despite the fact that this type of training has previously been shown to generalise well geographically [65], the shift in low-level features useful for the reconstruction pretext task was substantial between geographic regions. This is in contrast to previous work using M3LEO that found embeddings from DINO-based models with similar labels from different AOIs clustered tightly.
@@ 6 Acknowledgements
      Did you describe the limitations of your work? [Yes] See Appendix F.
@@
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix D.3
@@
      Did you mention the license of the assets? [Yes] See Appendix G.
@@ A.1 Chip counts
[h] Summary of component datasets in M3LEO, including number of chips and dataset size for each region (per year). Totals at the bottom are adjusted for multi-year datasets.
@@ A.2 Banding
Appendix B Marginal Distributions
The marginal distributions of ESAWC, MODISVEG, STRM mean elevation and GHSBUILTS can be seen in Figures B.1B.2B.3 and B.4 respectively.
Figure B.1: Marginal distribution for ESAWC. Note log scale.
Figure B.2: Marginal distribution for MODISVEG. Note log scale.
Figure B.3: Marginal distribution for SRTM. Mean elevation per-chip. Note log scale.
Figure B.4: Marginal distribution for GHSBUILTS. Note log scale.
Appendix C MAE Hyperparameters
Training hyperparameters for the MAE-based model can be seen in Table C.1. We followed [65], other than including additional AOIs in the pretraining set. A checkpoint for this model is available at huggingface.co/M3LEO.
Table C.1: Hyperparameter details for MAE pretraining used in main text.
MAE Pretraining
Encoder (Params) ViT-B [33] (88.8M)
Decoder (Params) Reconstruction [35] (5.5M)
Loss Function MSE
Input Image Size $448\times 448$
Output Image Size $448\times 448$
Patch Size $16\times 16$
Masking Type Random
Masking Ratio 0.75
Optimiser AdamW
Learning Rate 1.00E-04
No. Epochs 75
Input Dataset S1GRD
Channels Seasonal, VV, VH, VV-VH, Ascending
AOIs CONUS, Europe, China, Middle East, PAKIN, S. America
Appendix D Supervised Experiments
In addition to our explorations of distribution shift, we performed a small set of supervised experiments, reframing the auxiliary datasets as labelled tasks. S1GRD, GSSIC coherence and S2SRM were used as input datasets. For S1GRD, we trained models separately using the VV and VH bands only, and with the VV and VH bands stacked at the input to the model. For all S1GRD models, four seasonal summaries were used for each band, resulting in four input channels for the VV and VH models, and eight channels for the model using both VV and VH. For GSSIC coherence, the coherence band was used with a single date pair of delta 36 days taken for each season, resulting in four input channels. For the S2SRM models, we used the red, green and blue channels with one input channel from each of the months of March, June, September and December, totalling 12 input channels. We additionally trained models using both S2SRM RGB in combination with each of the other datasets separately, stacking the bands at the input to the model. All input data was taken from 2020. We upscaled low resolution input datasets to $448\times 448$ px before input to the model.
We excluded GUNW from use in our baseline experiments to avoid introducing the mixed availability of GUNW chips as a confounding factor.
ESAWC, AGB and GHSBUILTS labelled datasets were used as targets. ESAWC was formulated as semantic segmentation spanning 11 land use classes, for which segmentation accuracy is reported as mean intersection-over-union (mIoU). ESAWC data was used at the original resolution of $448\times 448$. Both AGB and GHSBUILTS are formulated as regression tasks (per-pixel). Results are reported using root mean square error (RMSE), in Mg ha-1 for AGB and in m2 built surface for GHSBUILTS. We resized labels for AGB and GHSBUILTS to a fixed size of $45\times 55$, to account for minor differences in dimension from chip-to-chip. We note that this means our pixels may not span exactly 1 hectare and therefore that results for AGB measured in Mg ha-1 are a heuristic rather than an absolute measure of biomass.
We excluded data from Europe in these experiments due to the absence of GUNW data in this region. This constraint was applied despite having not used GUNW in these baselines, as these experiments were completed prior to the exclusion of GUNW and repeating them was prohibitively expensive. All models were trained on the same random $10\%$ subset of the data. The spatial coverage of this subset is still similar to other popular large-scale EO datasets. We used the entire test set to compute the final metrics.
Models
We used a UNet-style architecture for all baseline experiments, following [103] with two major changes — halving the number of channels for all layers and replacing the up-convolutions with bilinear upsampling. For the AGB and GHSBUILTS regression tasks, we applied average pooling to the single-channel $448\times 448$ UNet output to achieve an output dimension of $45\times 55$. We opted not to use data augmentation. Selecting augmentations for SAR data is challenging — common choices such as rotation or flipping may introduce invariance to information specific to different instrument polarisations or orbital direction, for example. For further details on training and hyperparameters, see Appendix D.3.
D.1 Results
Results for all tasks are reported in Table D.1. For the experiments using a single data source as input, S1GRD using both polarisations (VV+VH) achieved the best result for the ESAWC (MIoU: 0.4185) and AGB (RMSE: 27.467 Mg ha-1) tasks. S2RGB achieved the best result for GHSBUILTS (RMSE: 131.968 m2). GSSIC coherence produced the worst result in all three cases (ESAWC MIoU: 0.2906, AGB RMSE: 39.314 Mg ha-1, GHSBUILTS RMSE: 131.968 m2). The best results for experiments using multiple data sources as input were achieved by fusion of S1GRD and S2RGB in all cases (ESAWC MIoU: 0.4634, AGB RMSE: 25.137 Mg ha-1, GHSBUILTS RMSE: 124.852 m2), with significant improvements over either S1GRD or S2RGB individually. Fusion of S2RGB with GSSIC improved results marginally (ESAWC MIoU: 0.4198, AGB RMSE: 28.550 Mg ha-1, GHSBUILTS RMSE: 131.300 m2) compared to either individual data source individually in all cases.
Table D.1: Baseline Evaluation Results for ESAWC, AGB, and GHSBUILTS tasks using our UNet models, outlined in Section D.
Input Dataset Bands ESAWC AGB GHSBUILTS
MIoU RMSE (Mg ha-1) RMSE (m2 built)
S1GRD VV 0.3976 28.573 152.259
S1GRD VH 0.3787 30.333 152.847
S1GRD VV+VH 0.4185 27.467 141.719
GSSIC Coherence 0.2906 39.314 196.242
S2RGB RGB 0.4094 28.689 131.968
S2RGB+S1GRD RGB+VV+VH 0.4634 25.137 124.852
S2RGB+GSSIC RGB+Coherence 0.4198 28.550 131.300
D.2 Discussion
In all experiments, S1GRD combining VV and VH polarisations performed similarly to S2RGB, although the gap was small. The inclusion of both polarisations for S1GRD increased performance uniformly compared to either individually. Unlike optical data, reflected SAR pulses are polarised differently depending on the terrain, and each measured polarisation contains unique information about the geometry of the imaged area. In all cases, performance was significantly improved by using both S1GRD and S2RGB data in fusion, despite giving similar performances individually, confirming a common finding in previous work.
Baseline experiments using GSSIC as the sole input data source performed significantly worse than for other input data types. This is likely explained in large part by the difference in resolution between GSSIC (90m, upsampled to 10m), S1GRD (10m) and S2RGB (10m). Performance improved slightly in all cases when GSSIC was included alongside in fusion with S2RGB. A nuanced approach is required for including coherence and interferometry, rather than naive direct input. Users may wish to use these data sources in self-supervised learning schemes — for example, pretraining models by constructing coherence or interferometry data from polarimetry pairs, or in a contrastive scheme alongside polarimetry data. Some work has been successful in using this type of pretraining [67]. Motivated by the success of polarimetry and coherence data in non-ML tasks [17, 18, 19, 20, 15, 13, 14, 16], we suggest that these data types generated from date-paired SAR acquisitions are likely to show stronger performance on change detection-type tasks, which we did not demonstrate here. M3LEO may provide useful pretraining data for these tasks.
D.3 Training Hyperparameters